A truck moving at 25 m/s south accelerates to 15 m/s south in 2.5 s. What is its acceleration?
1
Expert's answer
2020-11-03T10:31:53-0500

Solution

Acceleration is given by following equation of motion


"a=\frac{v-u}{t}"

Putting u and v and t value


"a=\frac{25-15}{2.5}=0.4" m/sec2